OpenStudy (anonymous):

What is the equation of the line perpendicular to –x + y = 7 and passing through (-1, -1)?

OpenStudy (anonymous):

-x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0 are the answer choices...can anyone help explain how i can solve this please?

OpenStudy (anonymous):

First, you need to determine the slope of the given equation.

OpenStudy (anonymous):

and the slope is (7,1) ?

OpenStudy (anonymous):

@Redcan ?

OpenStudy (anonymous):

\[y=x+7\] So comparing with the slope intercept form of the line, \[y=mx+b\] we see the slope is 1.

OpenStudy (anonymous):

ohhh ok right, so with a slope of 1, do i have to plug that into the answer choices?

OpenStudy (anonymous):

You want a line perpendicular to a line with slope one. So you need a line with new slope of -1/m.

OpenStudy (anonymous):

Draw a sketch and convince yourself.

OpenStudy (anonymous):

and perpendicular is creating a 90 degree angle right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Now use the point slope form of the line to find a line with slope -1 through (-1,-1).

OpenStudy (anonymous):

i dont think im drawing this correctly lol...ok so, on my graph i need to draw the first line, which is (-1,-1)

OpenStudy (anonymous):

and then the second line, has to create a 90 degree angle passing through the (-1,-1)?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[y-y_0=(x-x_0)m = (y-(-1))=(x-(-1))(-1)\]

OpenStudy (anonymous):

im trying to understand :( i just cant get it

OpenStudy (anonymous):

(-1,-1) is a point. y=mx+b is a line where m is slope.

OpenStudy (anonymous):

but i've gtg. sry

OpenStudy (mathmale):

@devmarie: Could you possibly explain what it is that you do understand here, as well as what you don't? What have you done so far towards solving this problem? Hint: Solve -x + y = 7 for y. Determine the slope of this straight line.